The length of a rectangle is 6 inches more than its width. When the lenght is doubled and the width is increased by 5 inches, the perimeter is 90 inches. What was the original length of the rectangle?
Answer by ankor@dixie-net.com(12701)

The length of a rectangle is 6 inches more than its width.
:
Let x = the width
Then
(x+6) = the length
:
When the length is doubled and the width is increased by 5 inches, the perimeter is 90 inches.
2(2(x+6)) + 2(x+5) = 90
:
Simplify divide thru by 2
2(x+6) + (x+5) = 45
2x + 12 + x + 5 = 45
3x + 17 = 45
3x = 45 - 17
3x = 28
x = 28/3
x = 9.333 inches is the width
:
What was the original length of the rectangle?
Length = 9.333 + 6 = 15.333 inches
:
:
Check using the perimeter statement:
2(30.666) + 2(14.333) =
61.333 + 28.666 = 89.999 ~ 90